Circular No. 6214

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA CASSIOPEIAE 1995 O. Ohshima, K. Ayani, M. Shimizu, and T. Yamamoto, Bisei As-

tronomical Observatory (BAO), Okayama, report that CCD spectrograms (range 390–700 nm, resolution 1.0 nm) of N Cas 1995 (IAUC 6213) were obtained on Aug. 26.736 and 27.5–27.7 UT with the BAO 1.0-m telescope. The spectra show very strong Hα emission (with weak or absent P-Cyg profile) and weak Fe II emission, and the strong continuum has absorption lines of Na I (D_{1} + D_{2}) and Ca II (H + K).

T. Iijima and L. Rosino, Asiago Astrophysical Observatory, write: “Optical spectra (range 410–510 nm, resolution 0.2 nm/pixel at 450 nm) of N Cas 1995 were obtained on Aug. 26.96 UT with the Asiago 1.2-m telescope (+ prismatic spectrograph). The spectra show a blue continuum and prominent, narrow emission lines of H I and Fe II multiplets 27, 37, 38, and 42. The FWHMs of H I lines are ∼ 300 km/s. H I and some Fe II lines have P-Cyg-type absorptions whose blueshifts are in the range from −280 to −400 km/s.”

M. Della Valle, Dipartimento di Astronomia dell’Universitá di Padova; and W. Marchiotto and G. Lercher, Institut für Astronomie, Universität Innsbruck, report: “Preliminary analysis of a spectrum of N Cas 1995, obtained under poor meteorological conditions at the Asiago Observatory’s Ekar Station with the 1.8-m telescope (+ Boller & Chivens spectrograph; range 470–680 nm) on Aug. 27.0 UT, confirms this object to be a galactic nova probably caught during its very early decline. The continuum is blue and dominated by emission lines of Balmer series and weaker lines due to Fe II (multiplets 42, 48, 74). The Na I line is detected in absorption. The expansion velocities inferred from the FWHMs of Hα and Hβ are ∼ 800 and 600 km/s, respectively.”

Ohshima also reports the following prediscovery unfiltered ST-6 CCD magnitudes obtained by K. Kosaka, Tamano, Okayama (comparison stars GSC 3668.1745, 3668.2028; SAO 21994): July 29.7 UT, [12.5; 30.7, 12.2:; 31.7, 11.6; Aug. 1.7, 11.0; 4.7, 10.4; 5.7, 9.9; 8.7, 9.6; 18.7, 9.8; 22.7, 9.5; 23.7, 9.2. G. M. Hurst, Basingstoke, England, reports prediscovery magnitudes from M. Collins, Everton, Bedfordshire (2415 films): Aug. 22.962 and 22.973, 9.0. Visual magnitude estimates: Aug. 27.177, 9.1 (J. E. Bortle, Stormville, NY); 27.79, 9.0 (L. Kiss, Szeged, Hungary); 28.118, 8.9 (P. Schmeer, Bischmisheim, Germany).

1995 August 28 Daniel W. E. Green